OpenStudy (anonymous):

HELP ME I'LL GIVE YOU A FREE CAT. Which number is between 0.5 and 5/8 ? A. 1/4 B. 8/28 C. 9/16 D. 19/ 28

OpenStudy (anonymous):

anyone ;c

OpenStudy (igreen):

Convert 5/8 into a decimal.

OpenStudy (igreen):

5 / 8 = ?

OpenStudy (anonymous):

What kind of cat?

OpenStudy (igreen):

Convert all the options to decimals too.

OpenStudy (anonymous):

.625??

OpenStudy (anonymous):

a cat.

OpenStudy (anonymous):

FOR MEEEEEEEEEEE?

OpenStudy (igreen):

Yep, 1/4 = ? 8/28 = ? 9/16 = ? 19/28 = ?

OpenStudy (anonymous):

totally. lol XD

OpenStudy (anonymous):

.25 0.285714 0.56 0.678571428571429 @igreen

OpenStudy (igreen):

Yep, so which number is between 0.5 and 0.625?

OpenStudy (anonymous):

is it A

OpenStudy (igreen):

No.. 0.25 is smaller than 0.5..so it's not in between 0.5 and 0.625

OpenStudy (igreen):

It has to be greater than 0.5 and less than 0.625

OpenStudy (igreen):

Just ignore the zero and the decimal point..

OpenStudy (igreen):

For all the options and those numbers..

OpenStudy (anonymous):

b

OpenStudy (igreen):

We're looking for a number between 50 and 62.5 We have: 25, 28.5, 56, and 67 Which number is it?

OpenStudy (anonymous):

ohhhhhhhhhh okay c

OpenStudy (igreen):

Yep..that's correct

OpenStudy (anonymous):

thank youuuuuuuuuuu

OpenStudy (yanasidlinskiy):

Not free cat, but GREEN cat ;) @iGreen

OpenStudy (igreen):

Yw

OpenStudy (anonymous):

i have 4 more questions btw XD

OpenStudy (igreen):

Oh, I get to keep @GreenCat as my personal slave? :D

OpenStudy (igreen):

Jk :3

OpenStudy (anonymous):

hahahah XD

OpenStudy (anonymous):

see yall lateraaa

OpenStudy (greencat):

oh no

OpenStudy (greencat):

it's a directioner!

OpenStudy (anonymous):

oh yeah 4 years going strong.

OpenStudy (igreen):

I don't listen to music..lol.

OpenStudy (anonymous):

.......... he is not human.

OpenStudy (anonymous):

well im a unicorn soo

OpenStudy (greencat):

it was a joke you were to think that i was saying oh no because ive been enslaved but it is because i see a directioner

OpenStudy (greencat):

igreen can have me.. as his owner. cuz im a cat and it's the cat that rules.

OpenStudy (greencat):

Congrats @iGreen for getting a free cat to own you.